OpenStudy (yuki):

Can someone explain me why the following limit is equal to 1 w/o using L'Hospital 's rule?

OpenStudy (yuki):

\[lin_{h \rightarrow 0} {\sin(h) \over h} = 1\]

OpenStudy (anonymous):

it's an identity.. and the only way to prove it is l'hopital.. i may be mistaken but thats the only way that i know how to prove that..

OpenStudy (amistre64):

as sin(h) gets really small; its about equal to "h"; and the smaller it gets the closer to 1 it gets

OpenStudy (amistre64):

type in sin(.00000001)/.00000001 and see what the calc gives yah

OpenStudy (yuki):

I get the intuitive idea, but is there any algebraic way?

OpenStudy (amistre64):

that is the algeraic way i think; it just is what it is :)

OpenStudy (amistre64):

you have an maybe if we express h in terms of radians? or is it already in terms of radians?

OpenStudy (anonymous):

you have to use something called the "Squeeze theorem"

OpenStudy (anonymous):

http://www.youtube.com/watch?v=Ve99biD1KtA

OpenStudy (anonymous):

@amistre the only way this is true is if h is in "radians". as a function, sine is a function of numbers, not of "degrees" or "radians". just like \[f(x)=x^2+1\] is a function of numbers. as such, it corresponds to the usual trig ratios of sides of triangles only if the angles are measured in radians

OpenStudy (anonymous):

and in fact, if you are using "degrees" as your input, the derivative of sine is NOT cosine. really.

OpenStudy (anonymous):

the basic geometric argument boils down to saying that \[\cos(x)<\frac{sin(x)}{x}<1\] near 0, and then as \[x->0\] the left hand side goes to 1 as well and therefore the limit is 1. but if you want a simple snap proof WITHOUT using derivatives there is not one.

OpenStudy (anonymous):

if you've had anlayss consider a sequence \[(x_n) \rightarrow 0 \] and use the continuity proof for \[(f(x_n)) \rightarrow f(x)\] and show that converges to 1 using a conventional \[\epsilon - \delta\] proof . If your more versed with topology you could define the filter converging to 0 and apply the function to it in a similar manner.

OpenStudy (anonymous):

with the filter showing it converges to1

OpenStudy (yuki):

Thanks a lot. That was one of the most clever proofs that I have ever seen. It was beautiful, too.